QUESTION #1: The base of a traingle sail is four times its height. If the area of the triangle is 162 square yards, find the base.
A = L x W 162 = 4x x=40.5
base = 4x
height = x
QUESTION #2: (3x-1)(9x2+3x+1)=0
3x-1=0 (3x-1)=0 (3x+1)=0
X=1/3 ? ?
Thank you!

Answer by checkley77(12844)

B=4H
THE AREA FORMULA FOR A TRIANGLE IS AREA=BH/2.
A=BH/2
162=4H*H/2
162=4H^2/2
4H^2=162*2
4H^2=324
H^2=324/4
H^2=81
H=SQRT81
H=9 YDS. ANS. FOR THE HEIGHT.
B=4*9=36 YDS. ANS. FOR THE BASE.
PROOF:
162=9*36/2
162=324/2
162=162
-------------------------
QUESTION #2: (3x-1)(9x2+3x+1)=0
3x-1=0 (3x-1)=0 (3x+1)=0
X=1/3 ? ?
(3x-1)(3x-1)(3x+1)=0
3x-1=0
3x=1
x=1/3 ans.
3x+1=0
3x=-1
x=-1/3 ans.